OpenStudy (anonymous):

In 1987, the Australian humpback whale population was 350 and has increased at a rate of about 14% each year since then. Write a function to model population growth. Use a graph to predict when the population will reach 20, 000.

OpenStudy (anonymous):

Use the compound interest formula you used in the other question.

OpenStudy (anonymous):

sooo wud it b 350(1+1.014)=20000

OpenStudy (anonymous):

Almost, the graph you will draw is 350(1+0.014)^n. Then you will go to the point where the graph equals 20,000 and see how many years it has been.

OpenStudy (anonymous):

290.9?

OpenStudy (anonymous):

290.99 is the correct answer, did you solve it using the graph?

OpenStudy (anonymous):

no im not sure how 2 graph it

OpenStudy (anonymous):

a)P(t) = 350(1.14)t; the population will reach 20,000 in about 30.9 yr. b)P(t) = 350(0.86)t; the population will reach 20,000 in about 19.5 yr. c)P(t) = 350(0.86)t; the population will reach 20,000 in about 39.1 yr.wwwwww d)P(t) = 350(1.14)t; the population will reach 20,000 in about 25.6 yr.

OpenStudy (anonymous):

Sorry, 290.99 is not the correct answer at all. The interest is 14% (which is 1.14), not 1.4% (which is 1.014). If you put 1.14 into the formula instead of 1.014 you should get the correct answer. Let me know how it goes.

OpenStudy (anonymous):

wud it b A?

OpenStudy (anonymous):

30.8?

OpenStudy (anonymous):

The answer on my calculator is 30.87543347. The correct rounding of this to one decimal place is 30.9. So the answer is A.